# Towards the Pedagogical Steering of Large Language Models for Tutoring: A Case Study with Modeling Productive Failure

### v1

@@ Abstract
One-to-one tutoring is one of the most efficient methods of teaching. Following the rise in popularity of Large Language Models (LLMs), there have been efforts to use them to create conversational tutoring systems, which can make the benefits of one-to-one tutoring accessible to everyone. However, current LLMs are primarily trained to be helpful assistants and thus lack crucial pedagogical skills. For example, they often quickly reveal the solution to the student and fail to plan for a richer multi-turn pedagogical interaction. To use LLMs in pedagogical scenarios, they need to be steered towards using effective teaching strategies: a problem we introduce as Pedagogical Steering and believe to be crucial for the efficient use of LLMs as tutors. We address this problem by formalizing a concept of tutoring strategy, and introducing StratL, an algorithm to model a strategy and use prompting to steer the LLM to follow this strategy. As a case study, we create a prototype tutor for high school math following Productive Failure (PF), an advanced and effective learning design. To validate our approach in a real-world setting, we run a field study with 17 high school students in Singapore. We quantitatively show that StratL succeeds in steering the LLM to follow a Productive Failure tutoring strategy. We also thoroughly investigate the existence of spillover effects on desirable properties of the LLM, like its ability to generate human-like answers. Based on these results, we highlight the challenges in Pedagogical Steering and suggest opportunities for further improvements. We further encourage follow-up research by releasing a dataset of Productive Failure problems and the code of our prototype and algorithm.
@@ 1 Introduction
One-to-one tutoring is an effective teaching method. Because of the efficiency of tutoring and its prohibitive cost, the challenge of finding affordable one-to-one tutoring methods is considered as one of the greatest problems in education: the so-called Two-Sigma Problem (Bloom 1984; Cohen, Kulik, and Kulik 1982). Following the rising popularity of Large Language Models (LLMs), there has been a growing interest in using them to make tutoring affordable by building conversational tutoring systems (Jurenka et al. 2024; Khan 2023) However, research empirically pointed out the lack of pedagogical properties of LLMs (Macina et al. 2023b; Tack and Piech 2022). These problems stem from the fact that most LLMs are instruction-tuned by optimizing for human preferences (Ouyang et al. 2022; Zhang et al. 2023). As a result, LLMs typically provide assistant-like responses. Yet a tutor’s goal is to maximize learning, not user satisfaction.
@@
These two goals can be contradictory. For example, if a student asks for help on a problem, directly providing them with the solution would likely maximize their satisfaction, and be the answer preferred by LLMs (Macina et al. 2023a). Yet, it is more efficient from a learning standpoint to promote students’ active engagement to let them cognitively engage with the problem, for example by providing an indirect hint or by asking a question (Freeman et al. 2014; Merrill et al. 1992). Moreover, existing LLM tutors are optimized and mostly evaluated for single-turn requests, while a tutoring dialog should be envisioned as a multi-turn teacher-student interaction where the teacher uses various pedagogical cues (Merrill et al. 1992). As LLMs are not natively made for tutoring, we introduce the fundamental shift from their user-serving goal to a pedagogically suitable one as the Pedagogical Steering problem.
@@
We present and evaluate StratL, an algorithm to address the Pedagogical Steering problem. To correct for a pedagogically suitable goal, we formally model the concept of a multi-turn tutoring strategy. We define a multi-turn strategy as a succession of single-turn pedagogical goals called tutoring intents. StratL dynamically redefines the tutoring intents after every student utterance using a transition graph and prompts the LLM to follow them for the next turn. We partner with learning scientists to integrate a strategy based on an advanced learning design known as Productive Failure (PF) (Kapur 2008). A meta-analysis of 53 studies showed that PF is more effective for human learning than other classical teaching methods (Sinha and Kapur 2021). PF can be seen as the most extreme version of not directly telling the answer to the students, as it intentionally lets them explore the problem and confront them with various suboptimal solutions, for example by using Socratic questioning. While PF works across learning subjects, this work focuses on high school math, as it has been most explored in PF research (Sinha and Kapur 2021), and prototypes ablations of StratL to evaluate it in real-world settings.
@@
We evaluate StratL and uncover practical limitations and insights on the Pedagogical Steering problem through two studies: a preliminary study with simulated students and a field study with 17 high school students in Singapore. Together, these studies show that StratL effectively steers the LLM to adhere to the Productive Failure tutoring strategy and does so with no negative spillover effects on other desirable properties of the LLM.
@@
  • We define the problem of Pedagogical Steering in the context of LLM multi-turn tutoring. We explain why instruction-tuned LLMs are misaligned for tutoring from a Learning Sciences standpoint, and review the current state of research from this perspective.
  • We introduce the StratL algorithm to model the concept of tutoring strategy and to prompt an LLM towards a given strategy using single-turn intents.
  • We run a field test in a Singaporean high school to evaluate the potential real-world impact and gather insights on the Pedagogical Steering problem. We measure the extent to which StratL and its ablations control the pedagogy followed by the LLM through Learning Sciences–inspired metrics and investigate the limitations.
  • We release a dataset of Productive Failure tutoring exercises. These exercises are particularly interesting because each has multiple suboptimal solutions and students often require multiple interactions to find the optimal one.
@@ 2 Background and Related Work
Pedagogical shortcomings of LLMs: Pedagogical Steering problem
@@
Recent advances enable state-of-the-art LLM-based systems to reliably perform middle– to high school–level mathematics, such as using prompting (Wu, Zhang, and Huang 2023; Kojima et al. 2022), multiple sampling strategies with consistency sampling (Wang et al. 2023) or refinement (Madaan et al. 2024; Shridhar et al. 2023; Li et al. 2023), fine-tuning (Zelikman et al. 2022; Yuan et al. 2023) or use of external tools like symbolic solvers or code interpreters (He-Yueya et al. 2023; Mialon et al. 2023). These reasoning capabilities could give LLMs the ability to serve as a basis for lesson-agnostic tutors.
Notwithstanding the promising capabilities of LLMs, several studies exhibited their poor pedagogical properties (Tack and Piech 2022). Among other flaws, most models often reveal the solution of the exercise to the students, even if prompted not to do so (Macina et al. 2023a). Their inefficiency as learning tools has also been measured more directly: when comparing students who were asked to solve problems using GPT-generated hints or human-generated ones, the human-generated hints group achieved substantially and statistically significantly higher learning gains (Pardos and Bhandari 2023).
@@
These pedagogical limitations stem from a core problem: LLMs are not optimized for tutoring. At a fundamental level, autoregressive LLMs are transformer-based models trained to take a sequence of tokens as input and predict the most likely following one (Zhao et al. 2023). After this initial pre-training process, an alignment step is necessary to turn these base models into useful assistants able to follow human instructions, like ChatGPT. The alignment is often performed by fine-tuning LLMs on human preferences (Zhang et al. 2023), mainly using a framework known as Reinforcement Learning with Human Feedback (RLHF) (Ouyang et al. 2022). During this process, LLMs learn to comply with requests by optimizing user satisfaction.
@@
LLM assistants thus generate texts that optimize for human preferences. This is misaligned with the way a tutor should behave. Indeed, when a student asks for help, the goal of the tutor should not necessarily be to maximize the student’s satisfaction, but rather their learning. For this, the tutor may have an intent (for example, asking back a question) different from what the student would like them to do (for example, being given the solution). This mismatch is one of the core limitations of LLMs as tutors.
LLMs for Tutoring
@@
Several works have attempted to address the lack of pedagogical properties of LLMs. The easiest and most direct approach to address this problem is to write an extensive prompt (Wang et al. 2024; Sonkar et al. 2023; Chowdhury, Zouhar, and Sachan 2024). However, enumerating all desired behaviors of a tutor in words is complex, especially for more nuanced pedagogical strategies.
@@
Another method considered is fine-tuning LLMs to efficient tutoring strategies. However, supervised fine-tuning is hampered by the scarcity of tutoring datasets, as their collection is costly and poses ethical concerns (Kasneci et al. 2023). Macina et al. (2023a) adopt an inverted Wizard-of-Oz approach by asking human teachers to tutor GPT-simulated students to create a comprehensive dataset of high-quality tutoring sessions and show that fine-tuning LLMs on this dataset improves their pedagogy. Similarly, the CLASS framework (Sonkar et al. 2023) uses an LLM to generate a synthetic conversational dataset with student-tutor interactions to fine-tune a smaller LLM. However, the quality of these conversations is not evaluated, and is prone to hallucinations. Moreover, fine-tuning is expensive and restricts the LLM to the specific teaching style present in the training data, thus requiring multiple datasets for different pedagogical approaches. The Bridge method (Wang et al. 2024), which we use as an inspiration for one of our baselines, uses an LLM to classify student errors and decide tutor interventions on the conversation history. Systems based on Bridge or CLASS let the LLM choose the tutoring strategy on its own, assuming its ability to make pedagogically sound choices. In contrast, our approach focused on letting learning scientists precisely control it. To do so, we employ a structured pedagogical intent selection using a transition graph based on classified student states, enabling experts to constrain the model to more faithfully follow a given pedagogy over multi-turn conversations.
Recent work reconsiders the CLASS framework from the point of view of alignment and tackles it using common alignment techniques such as RLHF (Sonkar et al. 2024). However, the data are synthetically generated which leads to questionable quality and bias, and as for fine-tuning, the need for data is the main drawback of this technique.
Productive Failure (PF)
@@
In education, “Instruction followed by Problem Solving” approaches, which directly instruct students on how to approach a problem and then let them practice, are widely used. However, Productive Failure (Kapur 2008, 2016) is a “Problem Solving followed by Instruction” learning design in which the learner attempts to solve open-ended problems before formally learning how to approach them, to build a deeper understanding of the topic. With PF, the teacher uses mostly questioning to let students cognitively engage with the problem. There is a large body of evidence that PF is effective with meta-analyses reporting an effect size of 0.87 standard deviations (Sinha and Kapur 2021), high by the learning sciences standards, and performing better than directly instructing students or telling them solutions.
@@
A successful problem-solving session in PF provides opportunities for the learner to explore the problem and build as many Representation and Solution Methods (RSMs) as possible, even if they fail at solving the problem (Kapur and Bielaczyc 2012). In simple terms, these RSMs can be thought of as mental representations of the problem, or partial solutions. The number of RSMs generated is one of the best predictors of how much students learn from PF (Kapur and Rummel 2012). Our tutoring strategy thus aims to promote the generation of such RSMs, and we will use their number as one of the evaluation metrics.
@@ 3.1 Formalism
We follow the formalization of the Dialog Tutoring task defined in (Macina et al. 2023b). We consider a tutoring dialog history $\mathcal{H}_{\leq n}=(u_{1},\ldots,u_{n})$ made of tutoring turns $u_{t}=(y_{t},x_{t})$ with $y_{t}$ the tutor’s utterance and $x_{t}$ the student’s utterance at turn $t$. $x_{t}\in\mathcal{V}^{\ast}$ and $y_{t}\in\mathcal{V}^{\ast}$ are finite sequences of tokens from a vocabulary $\mathcal{V}$. Each tutor utterance is associated with a set of tutoring intents $\mathbf{i}_{t}$ whose elements are chosen from a taxonomy of intents =$\mathcal{I}=\{I_{1},\ldots,I_{|\mathcal{I}|}\}$. An intent is defined as a goal of the tutor utterance, e.g. ‘correct the student’s mistake’. We denote by $K$ the grounding information, i.e. additional information provided to the LLM to give context. In our case, $K$ includes the tutoring exercise and its solution. The resulting generative conditional model is:
@@
We define a tutoring strategy predictor as a function $\phi$ that takes as input $(\mathcal{H}_{\leq t},K)$, the grounding and dialog history up to any turn $t$, and that outputs a set $\mathbf{i}_{t+1}\subseteq\mathcal{I}$ of single-turn tutoring intents for turn $t+1$. Abstractly, $\phi$ represents the decision process by which an expert tutor, given the context of the conversation, selects the best possible intents from a pedagogical standpoint to react to the student’s last utterance. The resulting conditional tutor model is:
@@
3.2 Tutoring Strategy – Learning Sciences Intuitions
Expert tutors base their interventions on observations about (1) the student’s emotional or knowledge state (Lepper et al. 1993), and (2) the state of the conversation and their previous interventions (Merrill et al. 1992). For example, a tutor should not react the same way to a computational mistake than to a reasoning one. Based on these intuitions, we suggest further refining the high-level intent selection function $\phi$, which abstractly depends on the dialog history and grounding, to a more practical function $\widetilde{\phi}$ that selects the next intents based on state features (1) and the previous intents (2). We propose the formalization developed in the next paragraph.
@@ 3.2 Tutoring Strategy – Learning Sciences Intuitions
Each turn is associated with a set $\mathbf{s}_{t}$ of state features describing the state of the student and of the conversation at turn $t$. The state features are chosen from a set $\mathcal{S}$ of possible features. A feature in $\mathcal{S}$ could, for example, be ‘the student is confused’ or ‘the student made a computational mistake’. We denote $\psi$ the abstract function that provides these features based on the dialog history and grounding. We approximate $\phi$ with a function $\widetilde{\phi}(\mathbf{i}_{n},\mathbf{s}_{n})=\mathbf{i}_{n+1}\subseteq\mathcal{I}$. $\widetilde{\phi}$ depends on the student state features $\mathbf{s}_{n}$ and previous tutor intents $\mathbf{i}_{n}$. Our final conditional generation model is:

### v2

@@ Abstract
One-to-one tutoring is one of the most efficient methods of teaching. With the growing popularity of Large Language Models (LLMs), there have been efforts to create LLM-based conversational tutors which can expand the benefits of one-to-one tutoring to everyone. However, current LLMs are trained primarily to be helpful assistants and lack crucial pedagogical skills. For example, they often quickly reveal the solution to the student and fail to plan for a richer multi-turn pedagogical interaction. To use LLMs in pedagogical settings, they need to be steered to use effective teaching strategies: a problem we introduce as Pedagogical Steering. We develop StratL, an algorithm to optimize LLM prompts and steer it to follow a predefined multi-turn tutoring plan represented as a transition graph. As a case study, we create a prototype tutor for high school math following Productive Failure (PF), an advanced and effective learning design. To validate our approach in a real-world setting, we run a field study with 17 high school students in Singapore and show that StratL succeeds in steering the LLM to follow the PF tutoring strategy. Finally, we highlight challenges in Pedagogical Steering of LLMs and offer opportunities for further improvements by publishing a dataset of PF problems and our code [^1] Dataset hyperlink; Code hyperlink.
Towards the Pedagogical Steering of Large Language Models for Tutoring: A Case Study with Modeling Productive Failure
Romain Puech1,2, Jakub Macina1,3, Julia Chatain4, Mrinmaya Sachan1, Manu Kapur4 1ETH Zurich, 2École polytechnique, 3ETH AI Center, 4Professorship for Learning Sciences and Higher Education, ETH Zurich, romain.puech@alumni.polytechnique.org
@@ 1 Introduction
One-to-one tutoring is an effective teaching method Bloom (1984). Because of the efficiency of tutoring and its prohibitive cost, the challenge of finding scalable and affordable one-to-one tutoring solutions is considered as one of the greatest problems in education: the so-called Two-Sigma Problem Bloom (1984); Cohen et al. (1982). Following the rising popularity of Large Language Models (LLMs), there has been a growing interest in using them to make tutoring affordable by building conversational tutoring systems Jurenka et al. (2024); Khan (2023). However, previous research has highlighted the lack of pedagogical properties of LLMs Macina et al. (2023b); Tack and Piech (2022). Most LLMs are instruction-tuned to optimize for a broad set of human preferences Ouyang et al. (2022); Zhang et al. (2023), leading to assistant-like responses. However, a tutor’s goal is to maximize learning, not user satisfaction.
@@
These two goals can be contradictory. For example, if a student asks for help, providing them directly with the solution would likely maximize their satisfaction and be the response preferred by LLMs Macina et al. (2023a). Yet, it is more efficient from a learning standpoint to promote students’ active engagement and let them cognitively engage with the problem, for example, by providing an indirect hint or asking a guiding question Freeman et al. (2014); Merrill et al. (1992). Moreover, existing LLM tutors are optimized and mostly evaluated for single-turn requests, while a tutoring dialog should be envisioned as a multi-turn teacher-student interaction where the teacher uses various pedagogical cues Merrill et al. (1992). As LLMs are not natively made for tutoring, we introduce the shift from their user-serving goal to a pedagogically suitable one as the Pedagogical Steering problem.
@@
We present and evaluate StratL, an algorithm to model a multi-turn tutoring strategy with LLMs. We define a multi-turn strategy as a succession of single-turn pedagogical goals called tutoring intents. StratL uses a transition graph to dynamically redefine the tutoring intents after every student utterance and prompts the LLM to follow these intents for the next turn. We partner with learning scientists to integrate a strategy based on an advanced learning design known as Productive Failure (PF) Kapur (2008, 2024), which is more effective for human learning than other classical teaching methods Sinha and Kapur (2021). PF can be seen as the most extreme version of not directly telling the answer to the students. In PF, students are encouraged to explore an ill-structured problem, for example by using Socratic questioning, and are confronted with various suboptimal solutions, even if it leads to failure. While PF works across learning subjects, this work focuses on high school math, as it has been most explored in PF research Sinha and Kapur (2021).
@@
We evaluate StratL and uncover practical limitations and insights on the Pedagogical Steering problem through a preliminary study with simulated students and a field study with 17 high school students in Singapore. Our results show that StratL effectively steers the LLM to adhere to the PF tutoring strategy and does so with no negative spillover effects on other desirable properties of the LLM.
@@
  • We define the problem of Pedagogical Steering in the context of LLM multi-turn tutoring. We introduce the StratL algorithm to model the concept of tutoring strategy and to prompt an LLM towards a given strategy using a sequence of single-turn intents.
  • We openly release the first dataset of PF tutoring exercises. These open-ended exercises have multiple suboptimal solutions and require multiple interactions to find the best one.
  • We run a field test in a Singaporean high school to evaluate the potential real-world impact of StratL. We measure the extent to which StratL and its ablations control the pedagogy through novel Learning Sciences–inspired metrics and investigate the limitations.
@@ 2 Background and Related Work
2.1 Pedagogical shortcomings of LLMs
@@
Recent advances enable state-of-the-art LLM-based systems to reliably perform middle– to high school–level mathematics, such as using prompting Wu et al. (2023); Kojima et al. (2022), multiple sampling strategies with consistency sampling Wang et al. (2023) or refinement Madaan et al. (2024); Shridhar et al. (2024); Li et al. (2023), fine-tuning Zelikman et al. (2022); Yuan et al. (2023) or use of external tools like symbolic solvers or code interpreters He-Yueya et al. (2023); Mialon et al. (2023).
@@
Notwithstanding the promising capabilities of LLMs, several studies exhibited their poor pedagogical properties Tack and Piech (2022). Among other flaws, most models often reveal the solution of the exercise to the students, even if prompted not to do so Macina et al. (2023a). These pedagogical limitations stem from a core problem: LLMs are not optimized for tutoring. After their initial pre-training, an alignment step is necessary to turn base language models into useful assistants capable of following human instructions, like ChatGPT. The alignment is often performed by fine-tuning LLMs on human preferences Zhang et al. (2023), mainly using a method known as Reinforcement Learning with Human Feedback (RLHF) Ouyang et al. (2022). During this process, LLMs learn to comply with requests by optimizing user satisfaction.
@@
LLM assistants thus generate texts that optimize for human preferences. This is misaligned with how a tutor should behave. Indeed, when a student asks for help, the tutor should not necessarily maximize the student’s satisfaction, but rather their learning. For this, the tutor may have an intent (for example, asking back a question) different from what the student would like them to do (for example, giving them the solution). This mismatch is one of the core limitations of LLMs as tutors.
2.2 LLMs for Tutoring
@@
Several works have attempted to address the lack of pedagogical properties of LLMs. The easiest and most direct approach is to write an extensive prompt Wang et al. (2024a); Sonkar et al. (2023); Chowdhury et al. (2024). However, enumerating all desired behaviors of a tutor in words is complex, especially for nuanced pedagogical strategies.
@@
Another method considered is fine-tuning LLMs to efficient tutoring strategies. However, supervised fine-tuning is hampered by the scarcity of tutoring datasets, as their collection is costly and poses ethical concerns Kasneci et al. (2023). Macina et al. (2023a) adopt an inverted Wizard-of-Oz approach by asking human teachers to tutor GPT-simulated students to create a comprehensive dataset of high-quality tutoring sessions and show that fine-tuning LLMs on this dataset improves their pedagogy. Similarly, the CLASS framework Sonkar et al. (2023) uses an LLM to generate a synthetic conversational dataset with student-tutor interactions to fine-tune a smaller LLM. However, the quality of these conversations is not evaluated, and is prone to hallucinations. Additionally, fine-tuning is expensive and restricts the LLM to the specific teaching style present in the training data, thus requiring multiple datasets for different pedagogical approaches. LearnLM Jurenka et al. (2024); LearnLM Team (2024) alleviates this issue by integrating pedagogical instruction following tuning, but still heavily relies on synthetic data and a costly co-training process where LearnLM is trained alongside a generic LLM. The Bridge method Wang et al. (2024a), which we use as an inspiration for one of our baselines, uses an LLM to classify student errors and decide tutor interventions on the conversation history. Systems based on Bridge let the LLM choose the tutoring strategy on its own, assuming its ability to make pedagogically sound choices.
In contrast, our approach focused on letting learning scientists control the tutoring behavior. To do so, we employ a structured pedagogical intent selection using a transition graph based on classified student states, enabling experts to constrain the model to more faithfully follow a given pedagogy over multi-turn conversations.
2.3 Productive Failure (PF)
@@
In education, “Instruction followed by Problem Solving” approaches, which directly instruct students on how to approach a problem and then let them practice, are widely used. However, PF Kapur (2008, 2016) rather uses the reversed order. PF is a “Problem Solving followed by Instruction” learning design, in which the learner attempts to solve open-ended problems before formally learning how to approach them, to build a deeper understanding of the topic. With PF, the teacher mostly uses questioning to let students engage cognitively with the problem. There is a large body of evidence that PF is effective with meta-analyses reporting an effect size of 0.87 standard deviations (high by the Learning Sciences standards) Sinha and Kapur (2021), when compared to traditional teaching.
@@
A successful problem-solving session in PF provides opportunities for the learner to explore the problem and build as many Representation and Solution Methods (RSMs) as possible, even if they fail at solving the problem Kapur and Bielaczyc (2012). In simple terms, these RSMs can be seen as mental representations of the problem, or unique solution attempts. The number of RSMs generated is one of the best predictors of how much students learn from PF Kapur and Rummel (2012). Our tutoring strategy thus aims to promote the generation of such RSMs, and we will use their number as one of the evaluation metrics.
@@ 3.1 Formalism
We extend the formalization of the Dialog Tutoring task defined in Macina et al. (2023b). We consider a tutoring dialog history $\mathcal{H}_{\leq n}=(u_{1},\ldots,u_{n})$ made of tutoring turns $u_{t}=(y_{t},x_{t})$ with $y_{t}$ the tutor’s utterance and $x_{t}$ the student’s utterance at turn $t$. Each tutor utterance is associated with a set of tutoring intents $\mathbf{i}_{t}$ whose elements are chosen from a taxonomy of intents =$\mathcal{I}=\{I_{1},\ldots,I_{|\mathcal{I}|}\}$. An intent is defined as a goal of the tutor utterance, e.g. ‘correct the student’s mistake’. We denote by $K$ the grounding information, i.e. additional information provided to the LLM to give context. In our case, $K$ includes the tutoring exercise and its solution. The resulting generative conditional model produces output:
@@
We define a tutoring strategy function as a function $\phi$ that takes as input $(K,\mathcal{H}_{\leq t})$, the grounding information and dialog history up to any turn $t$, and that outputs a set $\mathbf{i}_{t+1}\subseteq\mathcal{I}$ of single-turn tutoring intents for turn $t+1$. Abstractly, $\phi$ represents the decision process by which an expert tutor, given the context of the conversation, selects the best possible intents from a pedagogical standpoint to react to the student’s last utterance. This function will be replaced in practice by an approximation $\tilde{\phi}$ introduced in the next paragraph. The resulting conditional tutor model generates $y_{n+1}$ according to:
@@
Expert tutors base their interventions on observations about (1) the student’s emotional or knowledge state Lepper et al. (1993), and (2) the state of the conversation and their previous interventions Merrill et al. (1992). For example, a tutor does not react the same way to a computational mistake as to a reasoning one. Therefore, we use state features (1) we define in the next paragraph, and the previous tutoring intents (2) as inputs to our intent selection function $\tilde{\phi}$.
Each student turn is associated with a set $\mathbf{s}_{t}$ of state features describing the state of the student and of the conversation at turn $t$. The state features are chosen from a set $\mathcal{S}$ of possible features. A feature in $\mathcal{S}$ could, for example, be ‘the student is confused’ or ‘the student made a computational mistake’. We denote $\psi$ a classifier that provides these features based on the dialog history and grounding. Our intent selection function thus writes $\widetilde{\phi}(\mathbf{i}_{n},\mathbf{s}_{n})=\mathbf{i}_{n+1}\subseteq\mathcal{I}$. As previously noted, $\widetilde{\phi}$ depends on the student state features $\mathbf{s}_{n}$ and previous tutor intents $\mathbf{i}_{n}$. Our final conditional generation model produces output:
@@
Figure 2: Transition Graph of a PF Kapur (2008) Intent Selection. Nodes correspond to intents (listed in Table 5 of Appendix B.1). The variables used in the arrows’ transition conditions (‘a’, …, ‘m’) are codes for the state features (displayed in Figure 8 of Appendix B.2.1). At each turn, we transition from one set of intents to the next by following all satisfied arrows. Arrows with no starting node can be taken from any node. At turn 00, we start at node ‘Ask for the Next Step’. We describe the Learning Sciences justifications for this PF modeling in Appendix B.3.